(2x-6)^2 = -48
Solve for x.
Please help!

Answer by Alan3354(30993)

(2x-6)^2 = -48
Solve for x.
------------
(2x-6)^2 = -48
2x-6 = ±sqrt(-48) = ±4sqrt(-3)
x-3 = ±2sqrt(-3) = ±2isqrt(3)
x = 3 ±2i*sqrt(3) (i = sqrt(-1)